# Decadal analysis of sea surface temperature patterns, climatology, and anomalies in temperate coastal waters with Landsat-8 TIRS observations

### v1

@@ 3.1 Overview
In this study, we aim at analysing the spatiotemporal dynamics of SST in temperate coastal waters in South Australia with satellite-derived SST measurements. We first develop an operational approach to retrieving SST from Landsat-8 TIRS observations over large spatiotemporal scales, followed by introducing statistical analyses to examine the spatiotemporal patterns of SST in the study area. In Subsection 3.2, we briefly review the well-established equations underlying the radiative transfer process from the water-emitted radiance to the sensor-captured top-of-atmosphere radiance. In Subsection 3.3, we describe the derivation of atmospheric radiative variables via three radiative transfer simulations with the libRadtran software package [46]. We detail in Subsection 3.4 the calculation of atmospheric parameters from CAMS EAC4 atmosphere data which are required as inputs for the libRadtran simulations. Then, in Subsection 3.5, we calculate SST based on the radiative transfer equations and derived atmospheric radiative variables. In Subsection 3.6, we introduce a framework to parallelise the SST calculation for large spatiotemporal scales. Finally, we detail the methodologies for analysing the spatiotemporal patterns of SST, deriving the baseline SST climatology, and detecting anomalous SST events in Subsection 3.7.
@@ 3.2 Radiative Transfer Modelling
3.3 Derivation of Atmospheric Radiative/Transmission Variables
@@
In Eq. (1) the three variables $\tau(\lambda)$, $L_{u}(\lambda)$, and $L_{d}(\lambda)$ characterise the radiative/transmission properties of atmosphere. Following the approach in [14], three atmospheric simulations were conducted with the radiative transfer software package libRadtran [46], to derive these variables. It is worth noting that, these simulations were conducted under assumed atmospheric parameters for the purpose of solving the radiative transfer equation (Eq. (1)), rather than simulating with the actual atmospheric conditions.
@@
3.4 Calculation of Atmospheric Parameters
The parameters of atmospheric conditions (e.g., air pressure, temperature, and relative humidity) were required as inputs for the three libRadtran simulations in Subsection 3.3. Instead of applying a standard mid-latitude summer or winter atmosphere as the default atmospheric conditions, we chose to make use of spatiotemporally specific atmospheric conditions when and where the satellite observation was captured. Considering that accurate atmospheric corrections are critical in SST retrieval, location- and time-specific atmospheric conditions may help improve the quality of the retrieved SST values.
In this study, we used the CAMS EAC4 atmosphere data to derive spatiotemporally specific atmospheric conditions for each satellite observation. Considering that the spatial resolution of the CAMS EAC4 dataset is 0.75 degree longitude by 0.75 degree latitude and the temporal resolution is 3 hours, we divided the satellite image archive of interest into a spatiotemporal grid with each cell in the grid covering images over an area of 0.75 degree longitude by 0.75 degree latitude and a time period of 3 hours (as detailed later in Subsection 3.6).
For each cell, the corresponding mean sea level air pressure (single level value at sea surface, in Pa) was retrieved from the CAMS EAC4 data set. It was then converted from the single level air pressure at sea surface, to the vertical profile of air pressure at 60 discretised levels (from sea surface to top of atmosphere), based on the 60-level vertical discretisation model as defined in [47] and the converting coefficients for each model level as provided in CAMS EAC4:
$p^{(i)}=a^{(i)}+p_{0}\cdot b^{(i)}$ (8)
@@ 3.4 Calculation of Atmospheric Parameters
$CO_{2}^{(i)}=\left(1-q^{(i)}\right)\times 0.00053,$ (16)
In summary, for each cell that covers a spatial span of 0.75 degree longitude by 0.75 degree latitude and a temporal span of 3 hours, the spatiotemporally specific atmosphere variables, including air pressure $p^{(i)}$, air temperature $T^{(i)}$, air density $\rho^{(i)}$, water vapour content $q^{(i)}$, nitrogen concentration $N_{2}^{(i)}$, oxygen concentration $O_{2}^{(i)}$, and carbon dioxide concentration $CO_{2}^{(i)}$, ozone concentration $O_{3}^{(i)}$, and nitrogen dioxide concentration $NO_{2}^{(i)}$, at 60 model levels from sea surface to the top of atmosphere, were provided as inputs to the libRadtran simulations detailed in Subsection 3.3.
3.5 Calculation of SST
After determining the atmospheric radiative variables, i.e., $\tau(\lambda)$, $L_{u}(\lambda)$, and $L_{d}(\lambda)$, with the libRadtran simulations described in Subsection 3.3, we convolved these spectral quantities with the TIRS spectral response functions to convert them into broadband quantities that match with TIRS Band 10 (10.60–11.19 µm) and Band 11 (11.50–12.51 µm). We denote these broadband atmospheric radiative variables as $\tau^{\textnormal{B10}}$, $L_{u}^{\textnormal{B10}}$, and $L_{d}^{\textnormal{B10}}$ for Band 10, and $\tau^{\textnormal{B11}}$, $L_{u}^{\textnormal{B11}}$, and $L_{d}^{\textnormal{B11}}$ for Band 11. By convolving a lab measurement of the spectral emissivity of sea water with TIRS spectral response functions, the broadband emissivity of sea water at TIRS Bands 10 and 11 were determined to be $\varepsilon^{\textnormal{B10}}=0.9926$ and $\varepsilon^{\textnormal{B11}}=0.9877$, as documented in [14]. Then, deriving from Eq. (1), the black-body radiance of sea water at Band 10, $L_{s}^{\textnormal{B10}}$, can be calculated as:
3.6 Parallelisation for Large-Scale Processing
Given the spatiotemporal scale of this study, we divided the study area into multiple tiles to facilitate parallelised and efficient processing. Each tile encompasses a 0.75° × 0.75° longitude/latitude area, with its centre aligned to the sampling point of the CAMS EAC4 atmospheric data, as illustrated in Figure 6. Multitemporal TIRS images from 2014 to 2023 were extracted and processed independently for each tile. For efficiency, only the tiles intersecting the study area were processed, while water tiles outside the study area and pure-land tiles were skipped. These tiles were distributed across the workers of a Dask cluster on Amazon Web Services within CSIRO’s EASI platform. Please refer to the Data Availability section for details on data access.
3.7 Spatiotemporal Analyses
$T(d)=A\cdot\cos\left(\frac{2\pi d}{365}+\phi\right)+O,$ (19)
where $T(d)$ is the SST observation at the $d$th calendar day of a year, and the amplitude $A$, phase $\phi$, and offset $O$ are parameters to be determined through fitting. Figure 7 illustrates the procedure of extracting the baseline climatology for a randomly selected pixel. This approach was applied to every pixel within the study area to generate the climatology maps.
With the baseline climatology, we calculated the probability map of anomalous temperatures in the study area, which is defined as the percentage of SST observations from 2014 to 2023 that are greater than or less than 2°C from the baseline temperature. This threshold was empirically determined based on observations reported in the literature, which indicate that a temperature offset exceeding 2°C may cause adverse impacts on marine species, such as increasing the risk of mass coral mortality [49]. However, anomaly maps can be generated under other threshold values following the same approach described here. Based on the anomaly map, regions with high probabilities of anomalous SST were identified. To examine the seasonal distribution of SST anomalies, we further categorised the detected anomalous events by the warm months (January–March and October–December) and cool months (April–-September), and compared the likelihood of SST events between these two seasons.
4 Results and Discussions
4.1 Validation of SST Retrievals
The time-series plot in Figure 8a shows the comparison between satellite-derived SST data and in-situ SST measurements collected at the location of Port Lincoln buoy (135°54’01”E, 34°43’03”S). It was observed that the satellite-derived SST values aligned well with the in-situ measurements. It was also seen that the overall temperature trends, including seasonal variations and peaks, were captured by the satellite-derived SST. The scatter plot in Figure 8b presents the relationship between the satellite-derived and in-situ SST. The quantitative assessment shows a reasonably strong correlation between the two, with an $r^{2}$ value of 0.97 and a root mean square error (RMSE) of 0.38°C. The results for the Batemans Bay Offshore Buoy (150°20’13”E, 35°42’28”S), Coffs Harbour Offshore Buoy (153°15’32”E, 30°22’22”S), and Crowdy Head Offshore Buoy (152°51’08”E, 31°49’26”S) are shown in Figures 9a–b, 9c–d, and 9e–f, respectively. It was observed that the $r^{2}$ values are within the range of 0.85–0.90 and the RMSE values are lower than 0.75°C. The validation results in Figures 8 and 9 suggested that the SST retrievals obtained using the algorithm proposed in this study provide a reliable estimate of the temperature of sea surface waters.
4.2 Visualisation of SST Images
Figure 11 compares the SST images recorded by the MODIS Terra sensor (Figure 11a) and derived in this study from the Landsat-8 TIRS observation (Figure 11b). Both images were captured on June 24, 2020, with a time difference of $\sim$35 min, over waters surrounding the Jussieu Peninsula, South Australia. It was observed from the figure that both images shared a similar spatial pattern of SST, with a comparable range of SST values approximately between 10°C and 16°C. The MODIS Terra SST map (Figure 11a) displayed a relatively coarse spatial resolution (1000 m), with pixelated regions showing distinct blocks of temperature values. In contrast, the SST map derived from Landsat-8 TIRS (Figure 11b) showed a finer spatial resolution (100 m), displaying a smoother, more continuous temperature gradient, which highlighted smaller-scale thermal features that were less resovled in the MODIS data.
Figure 11: Comparison between the sea surface temperature (SST) images (a) recorded by the MODIS Terra sensor (1000 m resolution) and (b) derived in this study from Landsat-8 TIRS observations (100 m resolution).
4.3 Spatiotemporal SST Patterns
4.4 Daily Baseline Climatology
4.5 SST Anomalies
Figure 15 shows the probability of anomalous temperatures in the study area, calculated as the percentage of SST observations from 2014 to 2023 that are greater than or less than 2°C from the baseline climatology. The highest probabilities of anomalous temperatures were observed in the near-shore waters of Upper Spencer Gulf and Upper St Vincent Gulf, with the probability higher than 40% (Figure 15). The high rate of SST anomalies in these areas may be explained by the shallow depth of water. Shallow waters experience more rapid and extreme temperature fluctuations than deeper waters because their lower thermal inertia allows them to quickly heat up and cool down in response to sudden changes in solar radiation and atmospheric temperature, as documented in studies of the Gulf of California [53] and Long Island Sound, New York [54]. The restricted water circulation in Upper Spencer Gulf and Upper St Vincent Gulf may also play a role, as limited mixing with open ocean waters can lead to localised warming or cooling—a phenomenon also observed in a study of Hawaii coastal waters [52]. These combined factors make the near-shore, shallow areas in Upper Spencer Gulf and Upper St Vincent Gulf particularly prone to anomalous SST conditions.
4.6 Limitations and Prospects
In this study, we presented an operational method for retrieving SST from the TIRS sensor. We then introduced a novel algorithm for constructing a daily SST climatology, which served as a baseline for identifying anomalous SST events. These methods were applied to temperate coastal waters in South Australia over a ten-year period from 2014 to 2023. While the results demonstrated the potential of Landsat-8 TIRS in high-resolution SST monitoring, there are some limitations and several opportunities for future improvements.
Firstly, this study utilised the Landsat-8 TIRS sensor due to its relatively high spatial resolution (100 m) and over ten years of historical data. This enabled a detailed, long-term analysis of SST in South Australia’s coastal waters. Despite its advantages, Landsat-8 TIRS has a relatively large revisit time of 16 days, and cloud cover may potentially extend data gaps even further. Integrating SST information derived from other satellite sensors, such as Sentinel-3 SLSTR [25], MODIS [23], and VIIRS [24], could help mitigate data gaps and enhance the temporal resolution, although the lower spatial resolutions of these sensor (750 m–1 km) may present an additional challenge for identifying small-scale SST features. The more recently launched TIRS-2 instrument [56] onboard Landsat-9 could be leveraged as an additional source of SST data.
Secondly, this work focused on monitoring the thermal state of coastal waters. In order to provide an enhanced monitoring of South Australia’s coastal environment, it is worth exploring in future studies the synergisation of the thermal observations with information recorded by optical sensors, such as the multispectral instruments Landsat-8 OLI [50, 57], Sentinel-2 MSI [58, 59] and Sentinel-3 OLCI [58], and the hyperspectral instruments PACE [60], Tanager [61], EnMAP [62], and DESIS [63, 64]. In addition, sea surface salinity, which can be measured using spaceborne L-band radiometers [65], can also be integrated to enhance insights into the mixing dynamics of fresh and seawater.
Thirdly, this study focused on detecting historical SST anomaly events over the ten-year period from 2014–2023. These detected SST events provide important information on the spatiotemporal distribution of anomalies historically. Implementing real-time anomaly detection techniques, potentially using cloud-based processing platforms, could enable early warnings of significant SST deviations. Such advancements would be beneficial for fisheries, aquaculture, and environmental management by allowing timely responses to anomalous temperature fluctuations that may impact marine ecosystems.
5 Conclusion
This study demonstrated the capability of Landsat-8 TIRS in monitoring SST patterns, climatology, and anomalies in temperate coastal waters off the coast of South Australia over a ten-year period from 2014 to 2023. By developing an operational SST retrieval approach and establishing a novel high-resolution daily SST climatology, we examined the spatiotemporal patterns of SST and identified regions prone to anomalous temperature events. The high-resolution SST climatology derived in this study represents an improvement over existing datasets, offering enhanced spatial detail for detecting fine-scale temperature variations in coastal waters.
Data Availability
Acknowledgements
The authors sincerely thank Ms Gemma Kerrisk at CSIRO Environment for her assistance in in-situ data acquisition and preparation, and Dr Robert Woodcock and Mr Tisham Dhar at CSIRO Space and Astronomy for their support in data processing and indexing on the EASI ADS platform. The authors acknowledge the New South Wales Government Office of Environment and Heritage (OEH) for providing in-situ sea surface temperature data from the Manly Hydraulics Laboratory Waverider buoys.

### v2

@@ 3.1 Overview
In this study, we aim at analysing the spatiotemporal dynamics of SST in temperate coastal waters in South Australia with satellite-derived SST measurements. We first implement an operational approach to retrieving SST from Landsat-8 TIRS observations, followed by conducting statistical analyses to examine the spatiotemporal patterns, climatology, and anomalies of SST in the study area. In Subsection 3.2, we briefly review the well-established equations underlying the radiative transfer process from the water-emitted radiance to the sensor-captured top-of-atmosphere radiance. In Subsection 3.3, we describe the derivation of atmospheric radiative variables via three radiative transfer simulations with the libRadtran software package [47], and the calculation of atmospheric parameters from CAMS EAC4 atmosphere data which are required as inputs for the libRadtran simulations. Then, in Subsection 3.4, we calculate SST based on the radiative transfer equations and derived atmospheric radiative variables. In Subsection 3.5, we introduce a framework to parallelise the SST calculation for large spatiotemporal scales. Finally, we detail the methodologies for analysing the spatiotemporal patterns of SST, deriving the baseline SST climatology, and detecting anomalous SST events in Subsection 3.6.
@@ 3.2 Radiative Transfer Modelling
3.3 Derivation of Atmospheric Radiative Variables
@@
In Eq. (1), the three variables $\tau(\lambda)$, $L_{u}(\lambda)$, and $L_{d}(\lambda)$ characterise the radiative properties of atmosphere. Following the approach in [14], three atmospheric simulations were conducted with the radiative transfer software package libRadtran [47], to derive these variables. Details of the three libRadtran simulations are provided in Appendix A.
The parameters of atmospheric conditions (e.g., air pressure, temperature, and relative humidity) were required as inputs for the three libRadtran simulations. Instead of applying a standard mid-latitude summer or winter atmosphere as the default atmospheric conditions, we chose to make use of spatiotemporally specific atmospheric conditions when and where the satellite observation was captured. Considering that accurate atmospheric corrections are critical in SST retrieval, location- and time-specific atmospheric conditions may help improve the quality of the retrieved SST values.
In this study, we used the CAMS EAC4 atmosphere data to derive spatiotemporally specific atmospheric conditions for each satellite observation. Considering that the spatial resolution of the CAMS EAC4 dataset is 0.75 degree longitude by 0.75 degree latitude and the temporal resolution is 3 hours, we divided the TIRS image archive of interest into a spatiotemporal grid with each cell in the grid covering images over an area of 0.75 degree longitude by 0.75 degree latitude and a time period of 3 hours (as detailed later in Subsection 3.5). While this method improves the alignment between the coarse-resolution EAC4 atmospheric data and the fine-resolution TIRS data, some degree of mismatch remains for areas with complex atmospheric dynamics, suggesting a need for higher-resolution atmospheric datasets in future studies. For each cell that covers a spatial span of 0.75 degree longitude by 0.75 degree latitude and a temporal span of 3 hours, the spatiotemporally specific atmosphere variables, including air pressure $p^{(i)}$, air temperature $T^{(i)}$, air density $\rho^{(i)}$, water vapour content $q^{(i)}$, nitrogen concentration $N_{2}^{(i)}$, oxygen concentration $O_{2}^{(i)}$, and carbon dioxide concentration $CO_{2}^{(i)}$, ozone concentration $O_{3}^{(i)}$, and nitrogen dioxide concentration $NO_{2}^{(i)}$, at 60 model levels from sea surface to the top of atmosphere, were provided as inputs to the libRadtran simulations. Detailed calculations of these atmosphere variables are provided in Appendix B.
3.4 Calculation of SST
After determining the atmospheric radiative variables, i.e., $\tau(\lambda)$, $L_{u}(\lambda)$, and $L_{d}(\lambda)$, was described in Subsection 3.3, we convolved these spectral quantities with the TIRS spectral response functions to convert them into broadband quantities that match with TIRS Band 10 (10.60–11.19 µm) and Band 11 (11.50–12.51 µm). We denote these broadband atmospheric radiative variables as $\tau^{\textnormal{B10}}$, $L_{u}^{\textnormal{B10}}$, and $L_{d}^{\textnormal{B10}}$ for Band 10, and $\tau^{\textnormal{B11}}$, $L_{u}^{\textnormal{B11}}$, and $L_{d}^{\textnormal{B11}}$ for Band 11. By convolving a lab measurement of the spectral emissivity of sea water with TIRS spectral response functions, the broadband emissivity of sea water at TIRS Bands 10 and 11 were determined to be $\varepsilon^{\textnormal{B10}}=0.9926$ and $\varepsilon^{\textnormal{B11}}=0.9877$, as documented in [14]. Then, deriving from Eq. (1), the black-body radiance of sea water at Band 10, $L_{s}^{\textnormal{B10}}$, can be calculated as:
3.5 Parallelisation for Large-Scale Processing
Given the spatiotemporal scale of this study, we divided the study area into multiple tiles to facilitate parallelised and efficient processing. Each tile encompasses a 0.75° × 0.75° longitude/latitude area, with its centre aligned to the sampling point of the CAMS EAC4 atmospheric data, as illustrated in Figure 6. Multitemporal TIRS images from 2014 to 2023 were extracted and processed independently for each tile. For efficiency, only the tiles intersecting the study area were processed, while water tiles outside the study area and pure-land tiles were skipped. These tiles were distributed across the workers of a Dask cluster on Amazon Web Services within CSIRO’s EASI platform to accelerate data processing via parallelised computing.
3.6 Spatiotemporal Analyses
$T(d)=A\cdot\cos\left(\frac{2\pi d}{365}+\phi\right)+O,$ (4)
where $T(d)$ is the SST observation at the $d$th calendar day of a year. The selection of parameters, amplitude $A$, phase $\phi$, and offset $O$, was determined in this study through non-linear least-squares regression aiming at optimising the agreement between the fitted climatology and observed SST values. Figure 7 illustrates the procedure of extracting the baseline climatology for a randomly selected pixel. This approach was applied to every pixel within the study area to generate the climatology maps.
With the baseline climatology, we calculated the probability map of anomalous temperatures in the study area, which is defined as the percentage of SST observations from 2014 to 2023 that are greater than or less than 2°C from the baseline temperature. This threshold was empirically determined based on observations reported in the literature, which indicate that a temperature offset exceeding 2°C may cause adverse impacts on marine species, such as increasing the risk of mass coral mortality [49]. However, anomaly maps can be generated under other threshold values following the same approach described here. An anomaly map was generated for the study area by calculating the frequency of temperature anomalies, aggregating both positive and negative deviations from the climatological norm. Based on the anomaly map, regions with high probabilities of anomalous SST were identified. To examine the seasonal distribution of SST anomalies, we further categorised the detected anomalous events by the warm months (January–March and October–December) and cool months (April–-September), and compared the likelihood of SST events between these two seasons.
4 Results and Discussions
4.1 Validation of SST Retrievals
Figures 8 and 9 present the validation results of satellite-derived SST data against in-situ SST measurements. A total of 14, 57, 113, and 115 satellite/in-situ matchups were identified for the Port Lincoln Buoy, Batemans Bay Offshore Buoy, Coffs Harbour Offshore Buoy, and Crowdy Head Offshore Buoy, respectively. An in-situ SST measurement was considered as a valid match with a satellite-derived SST if the two were recorded within 100 m spatial distance and less than 30 minutes apart in time.
The time-series plot in Figure 8a shows the comparison between satellite-derived SST data and in-situ SST measurements collected at the location of Port Lincoln buoy (135°54’01”E, 34°43’03”S). It was observed that the satellite-derived SST values aligned well with the in-situ measurements. It was also seen that the overall temperature trends, including seasonal variations and peaks, were captured by the satellite-derived SST. The scatter plot in Figure 8b presents the relationship between the satellite-derived and in-situ SST. The quantitative assessment shows a reasonably strong correlation between the two, with an $r^{2}$ value of 0.97 and a root mean square error (RMSE) of 0.38°C.
The results for the Batemans Bay Offshore Buoy (150°20’13”E, 35°42’28”S), Coffs Harbour Offshore Buoy (153°15’32”E, 30°22’22”S), and Crowdy Head Offshore Buoy (152°51’08”E, 31°49’26”S) are shown in Figures 9a–b, 9c–d, and 9e–f, respectively. It was observed that the $r^{2}$ values are within the range of 0.85–0.90 and the RMSE values are lower than 0.75°C. The validation results in Figures 8 and 9 suggested that the SST retrievals obtained using the algorithm proposed in this study provide a reliable estimate of the temperature of sea surface waters.
4.2 Visualisation of SST Images
Figure 11 compares the SST images recorded by the MODIS Terra sensor (Figure 11a) and derived in this study from the Landsat-8 TIRS observation (Figure 11b). Both images were captured on June 24, 2020, with a time difference of $\sim$35 min, over waters surrounding the Jussieu Peninsula, South Australia. It was observed from the figure that both images shared a similar spatial pattern of SST, with a comparable range of SST values approximately between 10°C and 16°C. The MODIS Terra SST map (Figure 11a) displayed a relatively coarse spatial resolution (1 km), with pixelated regions showing distinct blocks of temperature values and less coherent temperature gradients. In contrast, the SST map derived from Landsat-8 TIRS (Figure 11b) showed a finer spatial resolution (100 m) than the MODIS data (Figure 11a), displaying spatially more coherent temperature gradients.
Figure 11: Comparison between the sea surface temperature (SST) images (a) recorded by the MODIS Terra sensor (1 km resolution) and (b) derived in this study from Landsat-8 TIRS observations (100 m resolution). The arrows indicate the direction and magnitude of SST gradients.
4.3 Spatiotemporal SST Patterns
4.4 Daily Baseline Climatology
4.5 SST Anomalies
Figure 15 shows the probability of anomalous temperatures in the study area, calculated as the percentage of SST observations from 2014 to 2023 that are greater than or less than 2°C from the baseline climatology. The highest probabilities of anomalous temperatures were observed in the near-shore waters of Upper Spencer Gulf and Upper St Vincent Gulf, with the probability higher than 40% (Figure 15). The high rate of SST anomalies in these areas may be explained by the shallow depth of water (see the bathymetry map of the study area in Figure 1). Shallow waters experience more rapid and extreme temperature fluctuations than deeper waters because their lower thermal inertia allows them to quickly heat up and cool down in response to sudden changes in solar radiation and atmospheric temperature, as documented in studies of the Gulf of California [53] and Long Island Sound, New York [54]. The restricted water circulation in Upper Spencer Gulf and Upper St Vincent Gulf may also play a role, as limited mixing with open ocean waters can lead to localised warming or cooling—a phenomenon also observed in a study of Hawaii coastal waters [52]. These combined factors make the near-shore, shallow areas in Upper Spencer Gulf and Upper St Vincent Gulf particularly prone to anomalous SST conditions.
4.6 Limitations and Prospects
In this study, we presented an operational method for retrieving SST from the TIRS sensor. We then introduced an algorithm for constructing daily SST climatology, which served as a baseline for identifying anomalous SST events. These methods were applied to temperate coastal waters in South Australia over a ten-year period from 2014 to 2023. While the results demonstrated the potential of Landsat-8 TIRS in high-resolution SST monitoring, there are some limitations and several opportunities for future improvements.
Firstly, this study utilised the Landsat-8 TIRS sensor due to its relatively high spatial resolution (100 m) and over ten years of historical data. This enabled a detailed, long-term analysis of SST in South Australia’s coastal waters. Despite its advantages, Landsat-8 TIRS has a relatively large revisit time of 16 days, and cloud cover may potentially extend data gaps even further. Integrating SST information derived from other satellite sensors, such as Sentinel-3 SLSTR [25], MODIS [23], and VIIRS [24], could help mitigate data gaps and enhance the temporal resolution, although the lower spatial resolutions of these sensor (750 m–1 km) may present an additional challenge for identifying small-scale SST features. The more recently launched TIRS-2 instrument [56] onboard Landsat-9 could be leveraged as an additional source of SST data. It is worth noting that integrating SST data from multiple satellite sensors presents several challenges, including spatial and temporal resolution mismatches, varying overpass times, and differences in sensor characteristics and atmospheric correction methods. Addressing these challenges would require approaches such as spatiotemporal harmonisation, cross-sensor calibration, and data fusion techniques.
Secondly, this work focused on monitoring the thermal state of coastal waters. In order to provide an enhanced monitoring of South Australia’s coastal environment, it is worth exploring in future studies the synergisation of the thermal observations with information recorded by optical sensors, such as the multispectral instruments Landsat-8 OLI [50, 57], Sentinel-2 MSI [58, 59] and Sentinel-3 OLCI [58], and the hyperspectral instruments PACE [60], Tanager [61], EnMAP [62], and DESIS [63, 64]. In addition, sea surface salinity, which can be measured using spaceborne L-band radiometers [65], can also be integrated to enhance insights into the mixing dynamics of fresh and seawater. For example, the remotely sensed salinity data were synergised with bio-optical observations to support water quality monitoring in Florida Bay, USA, as reported in [66].
Thirdly, this study focused on detecting historical SST anomaly events over the ten-year period from 2014–2023. These detected SST events provide important information on the spatiotemporal distribution of anomalies historically. Implementing real-time anomaly detection techniques, potentially using cloud-based processing platforms, could enable early warnings of significant SST deviations. Such advancements would be beneficial for fisheries, aquaculture, and environmental management by allowing timely responses to anomalous temperature fluctuations that may impact marine ecosystems. However, challenges remain regarding computational resource demands and data latency, which need to be addressed to ensure the reliability and timeliness of real-time anomaly detection systems.
5 Conclusion
This study demonstrated the capability of Landsat-8 TIRS in monitoring SST patterns, climatology, and anomalies in temperate coastal waters off the coast of South Australia over a ten-year period from 2014 to 2023. By developing an operational SST retrieval approach and establishing a high-resolution daily SST climatology, we examined the spatiotemporal patterns of SST and identified regions prone to anomalous temperature events. The high-resolution SST climatology derived in this study represents an improvement over existing datasets, offering enhanced spatial detail for detecting fine-scale temperature variations in coastal waters.
Data Availability
Acknowledgements
The authors would like to thank the anonymous reviewers for their important and insightful comments for improving this manuscript. The authors also thank Ms Gemma Kerrisk at CSIRO Environment for her assistance in in-situ data acquisition and preparation, and Dr Robert Woodcock and Mr Tisham Dhar at CSIRO Space and Astronomy for their support in data processing and indexing on the EASI ADS platform. The authors acknowledge the New South Wales Government Office of Environment and Heritage (OEH) for providing in-situ sea surface temperature data from the Manly Hydraulics Laboratory Waverider buoys.
A preprint version of this work is available at https://arxiv.org/abs/2503.05843.
Author Contributions
Yiqing Guo: Data curation, Formal analysis, Methodology, Visualization, Writing–original draft; Nagur Cherukuru: Conceptualization, Funding acquisition, Supervision, Writing–review & editing; Eric Lehmann: Conceptualization, Supervision, Writing–review & editing; Xiubin Qi: Data curation, Writing–review & editing; Mark Doubell: Funding acquisition, Writing–review & editing; S. L. Kesav Unnithan: Writing–review & editing; Ming Feng: Writing–review & editing.
Disclosure of interest
The authors declare no conflict of interest.
Appendix A Atmospheric simulations with libRadtran
Following the approach in [14], three atmospheric simulations were conducted with the radiative transfer software package libRadtran [47], to derive the atmospheric upwelling and downwelling radiances, $L_{u}(\lambda)$ and $L_{d}(\lambda)$, and the transmittance of atmosphere, $\tau(\lambda)$. It is worth noting that, these simulations were conducted under assumed atmospheric parameters. These assumed parameter values were empirically chosen for the purpose of solving the radiative transfer equation (Eq. (1)), rather than representing the actual atmospheric conditions.
@@ Appendix A Atmospheric simulations with libRadtran
Appendix B Calculation of Atmospheric Parameters
For each cell that covers a spatial span of 0.75 degree longitude by 0.75 degree latitude and a temporal span of 3 hours, the corresponding mean sea level air pressure (single level value at sea surface, in Pa) was retrieved from the CAMS EAC4 data set. It was then converted from the single level air pressure at sea surface, to the vertical profile of air pressure at 60 discretised levels (from sea surface to top of atmosphere), based on the 60-level vertical discretisation model as defined in [67] and the converting coefficients for each model level as provided in CAMS EAC4:
$p^{(i)}=a^{(i)}+p_{0}\cdot b^{(i)},$ (11)
@@ Appendix B Calculation of Atmospheric Parameters
$CO_{2}^{(i)}$ $=\left(1-q^{(i)}\right)\times 0.00053.$ (19)